OpenStudy (scorcher219396):

Can someone walk me through this problem? I understand the question and the math behind it, but have no idea where to start: How many tangent lines to the curve y=x/(x+1) pass through the point (1,2)? At which points do these tangent lines touch the curve?

jimthompson5910 (jim_thompson5910):

y = x/(x+1) is given

jimthompson5910 (jim_thompson5910):

what is y' ?

OpenStudy (scorcher219396):

y'=1/(x+1)^2

OpenStudy (scorcher219396):

Which is the slope of the curve, and the tangent line at that point

jimthompson5910 (jim_thompson5910):

Any point on the curve y = x/(x+1) is of the form (x,y) because y = x/(x+1), we can write that point as (x, x/(x+1) )

jimthompson5910 (jim_thompson5910):

We have these two points (1,2) which is given and what we want the tangent line to go through (x, x/(x+1) )

jimthompson5910 (jim_thompson5910):

Find the slope of the line through those two points to get